LOG-IN
Displaying report 1-1 of 1.
MIF (ITF Control)
Print this report.
kiwamu.izumi - 14:51, Tuesday 14 August 2018 (5801)Get code to link to this report
Temporary baffle PDs are good to go

The temprary baffle PDs were checked up today and confirmed to be functional.

Brief history:

The baffle PDs used at both end stations during phase 1 (39524008) had been taken out by the CRY group some months ago. Since the baffle PDs needed to be off from the structure to which the PDs had been glued on, they were removed from the structure. Later, the structures were black-coated to somewhat reduce the amount of scattered light. Anyhow, we now need to re-assemble the temporary baffle PDs back onto the structure. The thing is that we need to check up whether if the PDs are still functional before getting the stuff assembled.

Electronic test:

I looked at the voltage drop due to the diodes by using a multimeter in the PSL. The below is a list of the measured voltage drops.

  • Diode 1: 412 mV
  • DIode 2: 384 mV
  • Diode 3: 358 mV
  • Diode 4: 375 mV

Note that diode 3 was mechanically disconectted presumably due to somebody unintentionally having pulled the PD strongly. Also the attached is an illustration of this measurement.

Optical test:

Using the ALS setup on the PSL (5792), a 1064 nm light was prepared with an optical power of approximately 10 mW. The beam size was not measured this time. As the PDs were inserted to this beam path, they all consistently showed a DC photo current of about 2.5 mA. I used a voltmeter again. Even though this value is smaller by a factor of three than what it should be, I claim that this is good enough given that the beam size is not known.

Next things to do:

  • Glue the PDs back onto the structure.
  • Re-connect the PD that is currently disconnected.
Images attached to this report
5801_20180814075027_baffletest.png
Comments related to this report
yutaro.enomoto - 16:19, Tuesday 28 August 2018 (5915)

== Glueing of temporary baffle PDs for EXC ==
Ushiba, Enomoto

We press-and-connected the Kapton cables to a PD, which was accidentally unconnected, and glued 4 PDs on the plate that supports the PDs.
They are temporary ones and will be installed to EXC only for X arm commissioning.

See the first and second attached photos. The mounted PD unit shown here is now stored in EYC booth.
The glue we used is called EP-30.
We glued them at noon, and so we are planning to move it to EX area tomorrow morning.

We have two units of this temporary baffle PDs, and we did gluing for one unit today.
It is not planned that the other one is used anywhere, but we stored the unused unit in a rack just behind the entrance of PR3 clean booth (see the third and fourth pictures).

yutaro.enomoto - 11:24, Wednesday 29 August 2018 (5931)

The glued baffle PD unit were moved to EX area by Yokozawa Transportation Inc.
Now stored in EXA clean booth, as shown in the photo.